Mark Cotteleer owns a company that manufacturers canoes. Actual demand for Mark’s canoes during each season in 2012 through 2015 was as follow. Based on this data and the additive seasonal model, what will the demand level be for Mark’s canoes in the spring of 2017.

Year
Season 2012 2013 2014 2015
Winter 1400 1200 1000 900
Spring 1500 1400 1600 1500
Summer 1000 2100 2000 1900
Fall 600 750 650 500

You can use excel to solve this problem
Answer #1

Additive Model: yt = Tt + St + εt

For the four quarters the regression equation by the additive method is given as:

yt = β0 + β1t + β2S1 + β3S2 + β4S3 + β5S4 + εt

S1, S2, S3 and S4 are dummy variables for corresponding to seasons winter, spring, summer and fall respectively:

The combination of 0’s and 1’s for each of the dummy variables at each period indicate the quarter corresponding to the time series value.

Winter: S1 = 1, S2 = 0, S3 = 0, S4 = 0

Spring: S1 = 0, S2 = 1, S3 = 0 ,S4 = 0

Summer: S1 = 0, S2 = 0, S3 = 1, S4 = 0

Fall: S1 = 0, S2 = 0, S3 = 0, S4 = 0

Simplified regression equation is: yt = β0 + β1t + β2S1 + β3S2 + β4S3 + εt

Multiple regression is then done on with t, S1, S2, and S3 as the independent variables and the time series values yt as the dependent variable.

yt = 581.25 + 4.375t + 513.125S1 + 883.75S2 + 1129.375S3 + 0*S4

The excel model is as follows:

Formulas:

Solution:

Season

Forecast for year 2016

Winter

1168.75

Spring

1543.75

Summer

1793.75

Fall

668.75